# FedTAD: Topology-aware Data-free Knowledge Distillation for Subgraph Federated Learning

### v1

@@ 4.3 Topology-aware Knowledge Distillation
The proposed FedTAD servers as a post-processor following the vanilla model aggregators (e.g., FedAvg). Its objective is to transfer reliable knowledge from multi-client local models to further improve the aggregated model $\boldsymbol{\hat{w}}$. Specifically, in each round $t$, the server receives the local model weights $\{\boldsymbol{w}_{k}\}_{k\in S_{t}}$ and the class-wise knowledge reliability $\{\boldsymbol{\phi}_{k}\}_{k\in S_{t}}$ uploaded from the selected clients $S_{t}$.
@@
$\mathcal{L}_{diverg}=\sum_{k\in S_{t}}^{K}\sum_{c=1}^{C}\frac{\boldsymbol{\phi}_{k}^{c}}{\boldsymbol{\phi}^{c}}\sum_{v_{i}\in\mathcal{\hat{V}}_{c}}\text{KL}(f(v_{i},\hat{G};\boldsymbol{\hat{w}})||f(v_{i},\hat{G};\boldsymbol{w}_{k})),$ (10)
@@
17:           Update the global model $\boldsymbol{\hat{w}}$ via Eq. (11);
@@ 5.3 Experimental Analysis
Result 4: the answer to Q4. To answer Q4, we compare FedTAD with four baseline methods, including FedAvg, Fed-PUB, FedSage+, and FedGTA on Cora and ogbn-arxiv
@@
datasets with 20 participating clients. The activate fraction varies in $\{0.2,0.4,0.6,0.8,1.0\}$. The experimental results are presented in Table. 4. As observed, FedTAD consistently outperforms the baseline, indicating that FedTAD can maintain robustness with partial client participation.
@@
Figure 4: Experimental results for our ablation study.
@@ 6 Related Work
Graph Neural Networks. Earlier research on deep graph learning extends convolution to handle graphs Bruna et al. (2013) but comes with notable parameter counts. To this end, GCN Kipf and Welling (2016) simplifies graph convolution by utilizing a 1-order Chebyshev filter to capture local neighborhood information. Moreover, GAT Veličković et al. (2017) adopts graph attention, allowing weighted aggregation. Further details on GNN research can be found in comprehensive surveys  Wu et al. (2020); Zhou et al. (2020).
@@
Federated Graph Learning. Mainstream FGL studies can be divided into two settings: (i) Graph-FL: GCFL+ Xie et al. (2021) and FedStar Tan et al. (2023). Each client collects multiple independent graphs, aiming to collaboratively solve graph-level downstream tasks (e.g., graph classification). (ii) Subgraph-FL: each client holds a subgraph of an implicit global graph, aiming to solve node-level downstream tasks (e.g., node classification). Notably, there are currently two main challenges: (a) The global GNN performance degradation caused by subgraph heterogeneity (node and topology variation in multi-client subgraphs). Fed-PUB Baek et al. (2023) points out this heterogeneity stems from the various label distributions and addresses it through personalized technologies. FedGTA Li et al. (2023) utilizes a graph mixed moments to identify similar subgraphs, thereby achieving topology-aware aggregation. (b) The information loss caused by missing edges (connections implicitly exist between subgraphs across clients are lost due to distributed storage). Approaches for handling missing edges include FedSage+ Zhang et al. (2021) and FedGNN Wu et al. (2021).
@@
Data-Free Knowledge Distillation. This is a knowledge distillation technique that transfers the knowledge of the pre-trained teacher model to the student model through generated pseudo data instead of accessing the original data, including DeepImpression Nayak et al. (2019), DeepInversion Yin et al. (2020), and DFAL Chen et al. (2019). Recently, some studies introduced this strategy into FL to alleviate data heterogeneity, such as FedGen Zhu et al. (2021) and FedFTG Zhang et al. (2022). Unfortunately, they cannot be adapted to subgraph-FL scenarios as they can only generate Euclidean data rather than graph data.
@@ 7 Conclusion
In this paper, we explore the subgraph heterogeneity issue in subgraph-FL, providing an in-depth investigation of the individual impact of node and topology variation. We demonstrate that these variations result in differences in class-wise knowledge reliability. Building on this insight, we introduce FedTAD to correct the global model misled by unreliable knowledge by performing topology-aware data-free knowledge distillation. The experimental results demonstrate that FedTAD outperforms current leading approaches.

### v2

@@ 4.3 Topology-aware Knowledge Distillation
The proposed FedTAD servers as a post-processor following the vanilla model aggregators (e.g., FedAvg). Its objective is to transfer reliable knowledge from multi-client local models to further improve the aggregated model $\boldsymbol{\widetilde{w}}$. Specifically, in each round $t$, the server receives the local model weights $\{\boldsymbol{w}_{k}\}_{k\in S_{t}}$ and the class-wise knowledge reliability $\{\boldsymbol{\phi}_{k}\}_{k\in S_{t}}$ uploaded from the selected clients $S_{t}$.
@@
$\mathcal{L}_{diverg}=\sum_{k\in S_{t}}^{K}\sum_{c=1}^{C}\frac{\boldsymbol{\phi}_{k}^{c}}{\boldsymbol{\phi}^{c}}\sum_{v_{i}\in\mathcal{\hat{V}}_{c}}\text{KL}(f(v_{i},\hat{G};\boldsymbol{\widetilde{w}})||f(v_{i},\hat{G};\boldsymbol{w}_{k})),$ (10)
@@
17:           Update the global model $\boldsymbol{\widetilde{w}}$ via Eq. (11);
@@ 5.3 Experimental Analysis
Result 4: the answer to Q4. To answer Q4, we compare FedTAD with four baseline methods, including FedAvg, Fed-PUB, FedSage+, and FedGTA on Cora and ogbn-arxiv datasets with 20 participating clients. The activate fraction varies in $\{0.2,0.4,0.6,0.8,1.0\}$. The experimental results are presented in Table. 4. As observed, FedTAD consistently outperforms the baseline, indicating that FedTAD can maintain robustness with partial client participation.
@@
Figure 4: Experimental results for the ablation study.
@@ 6 Related Work
Graph Neural Networks.
Earlier research on deep graph learning extends convolution to handle graphs Bruna et al. (2013) but comes with notable parameter counts. To this end, GCN Kipf and Welling (2016) simplifies graph convolution by utilizing a 1-order Chebyshev filter to capture local neighborhood information. Moreover, GAT Veličković et al. (2017) adopts graph attention, allowing weighted aggregation. GraphSAGE Hamilton et al. (2017) introduces a variety of learnable aggregation functions for performing message aggregation. Further details on GNN research can be found in surveys  Wu et al. (2020); Zhou et al. (2020).
@@
Federated Graph Learning.
Mainstream FGL studies can be divided into two settings: (i) Graph-FL: GCFL+ Xie et al. (2021) and FedStar Tan et al. (2023). Each client collects multiple independent graphs, aiming to collaboratively solve graph-level downstream tasks (e.g., graph classification). (ii) Subgraph-FL: each client holds a subgraph of an implicit global graph, aiming to solve node-level downstream tasks (e.g., node classification). Notably, there are currently two main challenges: (a) The global GNN performance degradation caused by subgraph heterogeneity (node and topology variation in multi-client subgraphs). Fed-PUB Baek et al. (2023) points out this heterogeneity stems from the various label distributions and addresses it through personalized technologies. FedGTA Li et al. (2023) utilizes a graph mixed moments to identify similar subgraphs, thereby achieving topology-aware aggregation. (b) The information loss caused by missing edges (connections implicitly exist between subgraphs across clients are lost due to distributed storage). Approaches for handling missing edges include FedSage+ Zhang et al. (2021) and FedGNN Wu et al. (2021).
@@
Data-Free Knowledge Distillation.
This is a knowledge distillation technique that transfers the knowledge of the pre-trained teacher model to the student model through generated pseudo data instead of accessing the original data, including DeepImpression Nayak et al. (2019), DeepInversion Yin et al. (2020), and DFAL Chen et al. (2019). Recently, some studies introduced this strategy into FL to alleviate data heterogeneity, such as FedGen Zhu et al. (2021) and FedFTG Zhang et al. (2022). Unfortunately, they cannot be adapted to subgraph-FL scenarios as they can only generate Euclidean data rather than graph data.
@@ 7 Conclusion
In this paper, we investigate the subgraph heterogeneity problem in subgraph-FL. To the best of our knowledge, we are the first to provide an in-depth investigation of the subgraph heterogeneity problem by individually analyzing the impact of node and topology variation. We demonstrate that subgraph heterogeneity results in variation in the class-wise knowledge reliability. Building on this insight, we propose FedTAD, utilizing topology-aware data-free knowledge distillation to enable the local models to transfer their most reliable knowledge to correct the global model misled by unreliable knowledge. The experimental results demonstrate that FedTAD significantly outperforms state-of-the-art baselines.